A triangle has a perimeter of 105 inches. The lengths of the sides of the triangle are consecutive odd integers. Find the length of the sides and then use Heron’s Formula to find the area of a triangle. Show all work to receive credit
Answer by richwmiller(9144)

n+n+2+n+4=105
3n+6=105
3n=99
n=33
A=sqrt(s*(s-a)*(s-b)*(s-c))
where s=105/2=52.5
A=sqrt(52.5*(52.5-33)*(52.5-35)*(52.5-37))
A=526.965